OpenStudy (anonymous):

A school system is reducing the amount of dumpster loads of trash removed each week. In week 5, there were 45 dumpster loads of waste removed. In week 10, there were 30 dumpster loads removed. Assume that the reduction in the amount of waste each week is linear. Write an equation in function form to show the amount of trash removed each week.

OpenStudy (anonymous):

possible answers: f(x) = - 3x + 45 f(x) = 3x + 45 f(x) = - 3x + 60 f(x) = 3x + 60

undeadknight26 (undeadknight26):

you are in FLVS or OCVS?

OpenStudy (anonymous):

FLVS

undeadknight26 (undeadknight26):

I had this question...but i forgot how to solve it...

OpenStudy (anonymous):

Oh... Okay

undeadknight26 (undeadknight26):

let me see if i can find it...

undeadknight26 (undeadknight26):

lets see you have 2 points right?

undeadknight26 (undeadknight26):

5 over 45 and 10 over 30 right?

undeadknight26 (undeadknight26):

so 45 - 30/5 - 10

undeadknight26 (undeadknight26):

no...

OpenStudy (anonymous):

It didn't give me a graph or anything... Just the question

OpenStudy (anonymous):

Think of the data as two points on a graph, (5,45) and (10,30) find the slope, and then use point slope form to get your equation

undeadknight26 (undeadknight26):

what he said...that means i was on the right track...i should stop doubting myself lol...

OpenStudy (anonymous):

(45-30)/(5-10)=15/-5=-3 so slope =-3 point slope form using second point (10,30) y-30=-3(x-10) y-30=-3x+30 y=-3x+60

OpenStudy (anonymous):

Okay so then the answer would be D?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

Oh duh! The negative!